Given μ = -10, σ = 5, n=500, find 𝜇𝑋̅ and 𝜎 2 𝑋̅ Solutio

1
Expert's answer
2022-05-16T17:41:37-0400
"\mu_{\bar{X}}=\mu=-10"

"\sigma_{\bar{X}}^2=\dfrac{\sigma^2}{n}=\dfrac{5^2}{500}=0.05"